OpenStudy (tennistar):

Simplify: the square root of c (3 + 2 the square root of 5)

OpenStudy (campbell_st):

can you use the draw option to write the expression

OpenStudy (tennistar):



OpenStudy (campbell_st):

expand so its \[\sqrt{c} \times 3 + 2\times \sqrt{5}\times \sqrt{c}\] this can be simplified\[3\sqrt{c} + 2\sqrt{5\times c}\] which gives the final answer of \[3\sqrt{c} + 2\sqrt{5c}\]

OpenStudy (tennistar):

Are you sure this is correct?

OpenStudy (tennistar):

I cant mess up on this question

OpenStudy (campbell_st):

yes... provided what you wrote is correct.

OpenStudy (campbell_st):

ask someone else if you are unsure

OpenStudy (tennistar):

the real question